OpenStudy (anonymous):

Classify the real number in as many ways as possible 65/9

OpenStudy (anonymous):

Here is a guide with examples: http://www.ck12.org/na/Classifying-Numbers-1/lesson/user:Dysart/Classifying-Numbers/ \( \Large ☺\)

OpenStudy (anonymous):

i still dont understand

OpenStudy (michele_laino):

1)is a rational number since it is the ratio between two integers, namely 65 and 9 2) it is an improper fraction, since the denominator is greater than the denominator, namely 65>9 3) it is a mixed number since I can rewrite it as follows: 7 2/9

OpenStudy (michele_laino):

please note that, we have: 65:9 = 7 and the remainder is 2

OpenStudy (anonymous):

@Michele_Laino It is also a positive and real number. and 2 and 3 basically mean the same thing. I don't mean to annoy, but I was wondering myself where the limit of this classification ends. So I googled this form of question and the example I found limits it to: real, rational, irrational, integer, whole and natural. If such limitation exists then I'd say 'real' and 'rational' is the best classification. @Ariarose Are you familiar with those group of numbers?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

http://www.mathsisfun.com/sets/number-types.html Try to see to what groups your number belong

OpenStudy (anonymous):

i dont understand this at all...im sorry..

OpenStudy (anonymous):

Alright, let's see. Is it a whole number?

OpenStudy (anonymous):

no....? because if you divide it it would equal 7.222 i think

OpenStudy (anonymous):

Alright. So it isn't. As Michele has mentioned you can also see that the division ends up with a remainder. Either way it is not a whole number. So the it doesn't belong in the first two groups: It isn't a Natural number or an Integer. Next are rational numbers. Rational numbers are any numbers that can be described by division of two integers. That might sound confusing at first, but here are some examples: $$ \frac{1}{2}, \frac{15}{3}, 5, \frac{-2}{3} $$You can see that \(5\) is also rational because it is also \(\frac{5}{1}\). So all whole numbers are also rational. So, is this number a rational number?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Correct. So if it is a Rational number it is not an Irrational number, which is simply the opposite.. The example I found does not deal with Complex numbers so I'll leave it aside. The normal numbers that you know: \(4, \frac{1}{2}, -3, \pi...\) are all real numbers, and this is not an exception. So it is a Real number and a Rational number. Makes sense?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ok, good. Now look at the example again, they use more numbers there.

OpenStudy (anonymous):

ok